OpenStudy (anonymous):

find the x and y intercepts 7x-4y=56

OpenStudy (anonymous):

x-intercept = 8 (put y=0) y-intercept = -14 (put x=0)

OpenStudy (anonymous):

x-intercept (put y=0) 7x=56 x=56/7=8 y-intercept (put x=0) -4y=56 y=56/-4=-14